One of the two digits of a two-digit number is three times the other digit. If you interchange the digits of this two-digit number and add the resulting number to the original number, you get 88. What is the original number? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the digit at tens place be x then digit at ones place will be 3x.
Original two-digit number = 10x + 3x
After interchanging the digits, the new number = 30x + x
According to the question,
(30x+x)+(10x+3x)=88
⇒31x+13x=88
⇒44x=88
⇒x=2
Originalnumber=10x+3x=13x=13×2=26

Thanks